OpenStudy (anonymous):

Verify the following. 1. 1-cot^2 x sin^2 x = sin^2 x NOTE: x= the zero degree sign I

OpenStudy (anonymous):

I wasn't able to write the correct sign on here so i used x instead

OpenStudy (anonymous):

theta?

OpenStudy (anonymous):

yes that sign thank you

OpenStudy (anonymous):

cotan and sin are multiplied?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[1-\cot ^{2} \theta \sin ^{2} \theta = \sin ^{2}\theta\]

OpenStudy (anonymous):

^ yes thats how the problem looks

OpenStudy (anonymous):

maybe divide through by sin^2 theta

OpenStudy (anonymous):

which will give you 1/ sin ^2 x - cot^2 x on the left and 1 on the right

OpenStudy (anonymous):

cot = 1/tan, tan = sin/cos, so cot = cos/tan

OpenStudy (anonymous):

sorry not cos/tan, cos/sin

OpenStudy (anonymous):

I think the problem might want you to use a trig identity to prove \[\sin ^{2}\]

OpenStudy (anonymous):

yes it wants me to use the identities

OpenStudy (anonymous):

\[\sin ^{2}\theta + \cos ^{2}\theta = 1\]

OpenStudy (anonymous):

yes thats the one.

OpenStudy (anonymous):

so multiply through by sin^2 x. then change cot^2 x into (cos^2 x)/(sin^2 x) so you now have...

OpenStudy (anonymous):

\[1 + \cot ^{2}\theta = \csc ^{2}\theta \]

OpenStudy (anonymous):

\[\frac{1 - \cos^{2} \theta} {\sin^2 \theta} = 1\]

OpenStudy (anonymous):

then just multiply by sin^2 theta and you have the identity

OpenStudy (anonymous):

now back to call of duty :D

OpenStudy (anonymous):

OK thank you both so much now i just have to follow through to make sure that i understand it

OpenStudy (anonymous):

try and do the steps i did, and tell me if you get stuck

OpenStudy (anonymous):

instead of multiplying sin though she needs to divide

OpenStudy (anonymous):

your identity is equal to sin not cosecant

OpenStudy (anonymous):

yes thats right, at the start divide by sin^2

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

done it?

OpenStudy (anonymous):

still working on it give me a minute

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

ok in stuck and confused

OpenStudy (anonymous):

show me where your up to

OpenStudy (anonymous):

you have to verify that the equation equals sin^2 theta

OpenStudy (anonymous):

your trying to prove that the left side = the right side.

OpenStudy (anonymous):

so you would just look at the problem being 1-cot^2thetasin^2theta

OpenStudy (anonymous):

so if you can get the equation into the form \[1 = \cos^{2} + \sin^{2}\] Which we know is always true. then you have proved it

OpenStudy (anonymous):

your trying to manipulate it into that form

OpenStudy (anonymous):

or one of these http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

here were trying to get it into the "pythagorean identity"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so youve divided by sin^2 x

OpenStudy (anonymous):

\[\frac{1}{\sin^{2} x} - \cot^{2} x = 1\] and you got that?

OpenStudy (anonymous):

wait when i divide sin^2 x i divide it on both sides correct?

OpenStudy (anonymous):

then using the identity that \[\cot^2 x = \frac{1}{\tan^2 x} = \frac{\cos^2 x}{\sin^2 x}\]

OpenStudy (anonymous):

yeah on both sides, to keep it balanced

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

but if i divide it on both sides i would get 1-cot^2x=1 i dont get how you got something different

OpenStudy (anonymous):

okay so because \[\cot^2 x = \frac{1}{\tan^2 x} = \frac{\cos^2 x}{\sin^2 x}\] we can now rewrite it as... \[\frac{1}{\sin^{2} x} - \frac{\cos^2 x}{\sin^2 x}\ = 1\] so we replaced the cot term with the cos/sin, is there anything you dont get?

OpenStudy (anonymous):

no you have to divide the 1 at the front by sin^2 too

OpenStudy (anonymous):

got to do the same thing to every term

OpenStudy (anonymous):

to keep it nice and balanced

OpenStudy (anonymous):

do you know that sin = opposite / hypotenuse, cos = adjacent/ hypotenuse, tan = opposite/adjacent ?

OpenStudy (anonymous):

yes now i understand how you got to that point

OpenStudy (anonymous):

so now that i get to the point 1/sin^2x-cos^2x/sin^2x can i join them together and write it as 1-cos^2x/sin^2x which will then equal sin^2/sin^2= 1 ?

OpenStudy (anonymous):

you need to use parenthesis

OpenStudy (anonymous):

no keep it like this \[\frac{1}{\sin^{2} x} - \cot^{2} x = 1\] have you got to that stage?

OpenStudy (anonymous):

what you should have jusging by what you've written it should be csc^2(x)-cot^2(x)

OpenStudy (anonymous):

im solving it for sin^2 x + cos^2 x = 1

OpenStudy (anonymous):

and simplify jame's equation and you have your identity

OpenStudy (anonymous):

you take over veritas

OpenStudy (anonymous):

\[\csc ^{2}\theta + \cot ^{2}\theta= 1 \rightarrow 1- \cot ^{2}\theta = \csc ^{2}\theta \]

OpenStudy (anonymous):

ok i am very confused right now so now that i got to 1/sin^2 x - cot^2x

OpenStudy (anonymous):

they dont have the same denominator so do i still change the cot^2 x into the csc^2 x i am very confused with that part

OpenStudy (anonymous):

wait no now i get it

OpenStudy (anonymous):

you changed the 1/sin^2 into the csc right?

OpenStudy (anonymous):

yes he did

OpenStudy (anonymous):

exactly that, have you solved it now? :)

OpenStudy (anonymous):

no i just got csc^2x-cot^2x=1 do i add the cot^2x over to the other side?

OpenStudy (anonymous):

yeah cot to the other side and you have nailed it

OpenStudy (anonymous):

ok thank you both for your patience and for your help i really appreciate it

OpenStudy (anonymous):

no problem, glad you understand it

OpenStudy (anonymous):

Me too lol

OpenStudy (anonymous):

Just anoying that it doesnt get much easier. Its a constant struggle to grasp the next thing im learning :(

OpenStudy (anonymous):

i definitely understand where your coming from ='( lol